# Sharp Tools: How Developers Wield Agentic AI in Real Software Engineering Tasks

### v2

@@ Abstract
Software Engineering Agents (SWE agents) can autonomously perform development tasks on benchmarks like SWE Bench, but still face challenges when tackling complex and ambiguous real-world tasks. Consequently, SWE agents are often designed to allow interactivity with developers, enabling collaborative problem-solving. To understand how developers collaborate with SWE agents and the communication challenges that arise in such interactions, we observed 19 developers using an in-IDE agent to resolve 33 open issues in repositories to which they had previously contributed. Participants successfully resolved about half of these issues, with participants solving issues incrementally having greater success than those using a one-shot approach. Participants who actively collaborated with the agent and iterated on its outputs were also more successful, though they faced challenges in trusting the agent’s responses and collaborating on debugging and testing. These results have implications for successful developer-agent collaborations, and for the design of more effective SWE agents.
Index Terms:
Agent-based systems, software development tools, collaboration, communication, developer experience
@@ I Introduction
SWE agents have shown promising results on established benchmarks of software development tasks, such as fully autonomously solving many of the GitHub issues in SWE-Bench [3]. However, they still struggle to solve some issues due to several factors, such as the complexity and ambiguity of the issues, lack of tacit knowledge, and limited access to the full development environment [4, 5]. To overcome these limitations, many SWE agents are designed to interact with developers, enabling a collaborative approach to problem-solving. This interaction allows developers and agents to share responsibilities such as understanding the codebase, localising, coding, debugging, testing, and reviewing each other’s work.
Figure 1 shows an example of one such SWE agent – Cursor Agent in the Cursor integrated development environment (IDE) – responding to a developer’s prompt that has been copied from a GitHub bug report [6]. In the top part of the figure, the agent searches the codebase, describes the issue based on this search, and tries to implements a fix. After prompting, the agent then adds a unit test and runs it.
While reviewing the agent’s changes, the developer notices a functional call erroneously changed by the agent (Figure 2, bottom), and rejects that part of the change, before reviewing the remaining 5 changes.
There is reason to believe that collaborations between agents and developers will not be smooth and painless. For non-agentic tools, prior research has found that developers sometimes struggle with understanding and trusting the AI’s output, hindering effective collaboration [7, 8, 9]. Is this also the case for SWE agents? On one hand, unlike non-agentic tools that generate a single response and act as a black box, agents can provide developers with information about their process by showing their work as they go. Further, their capability to make decisions autonomously allows for developers to express higher-level goals rather than low-level instructions. On the other hand, the broader scope of agents’ decisions and subsequent actions can exacerbate trust issues and place a higher cognitive load on the developer of understanding both the agent’s reasoning trajectory as well as mutations it performs to the codebase. The answer is yet unclear, due to a lack of research on real-world usage of SWE agents [2].
Figure 1: Example usage of the Cursor Agent
Figure 2: A suggested code edit diff by the agent.
@@ 2. RQ2:
    What are the communication barriers to collaborations between software developers and SWE agents when addressing open issues in active codebases?
@@ 3. RQ3:
    What factors influence participants’ success when collaborating with a SWE agent to solve open issues in active codebases?
@@ II-A Agentic Tools for Software Development
In software development, there has been a rapid rise in the number of agentic tools in research and in practice [10]. Many of these tools aim to solve end-to-end software engineering tasks, as reflected in the evolution of benchmarks, such as SWE-bench [3], a set of real previously-solved GitHub issues, and LiveCodeBench [13], which adds new issues over time. Similarly, SWE-Lancer [4] uses real-world freelance software engineering tasks to evaluate AI tools. While these benchmarks aim to evaluate real-world performance, they all test the autonomous performance of agentic tools, rather than their performance with human-in-the-loop. As Liu and colleagues report, most research-based agentic systems target maximum automation, where human involvement is limited to specification of the problem statement [2].
@@ II-B Developer Experiences with AI Tools
Prior research has also explored the strategies that developers apply when using AI [17], especially prompt engineering strategies for developers [18, 19, 20]. Barke and colleagues found that developers used GitHub Copilot code completions and chat in different ways at different stages of solving an issue [8]. Prior work has also found that several development tasks are time-consuming or challenging when working with AI developer tools, including communicating the required context to AI tools [8, 17], verifying AI-generated code recommendations [21], and debugging incorrect AI-generated code [9]. Developer studies are mixed about whether AI increases developer efficiency [9] or decreases it [22]. Studies have also found that developers often use AI tools as learning aids and information sources [9, 23], and have concerns about the quality and reliability of generated code [21, 7], including code security  [9, 22]. While these studies have richly explored developer interactions with AI code completion [21, 9, 8, 24] and AI-based chat [25, 26, 27, 28], there is a dearth of developer experience studies on the use of SWE agents. Yet, SWE agents offer a unique interaction paradigm, differing from non-agentic tools in their capability to perform complex tasks autonomously and take feedback from the environment [2].
@@ III Study Design
To answer our research questions, we aimed to observe developers collaborating with a SWE agent in a highly ecologically valid setting. To achieve this, we asked 19 professional software developers to use the Cursor Agent to resolve issues in a codebase to which they have previously contributed.
@@ III-A Tool Selection
In selecting the appropriate SWE agent for our study, we considered the spectrum of human-AI collaboration capabilities, ranging from highly autonomous agents to those that enable robust interaction with developers. We reviewed several agents, starting with fully autonomous options like AutoCodeRover [35], and moving to agents that allow a moderate level of human-AI collaboration. These moderate collaboration tools often operate outside the IDE and primarily rely on user interaction through prompting, such as Devin [36], Aider [37],OpenHands [38] and GitHub Copilot Workspaces [39]. Finally, we considered agents that operate within an IDE, facilitating high collaboration not only through prompting but also by providing developers access to the full development environment. Examples include Cursor Agent [16], VSCode Agent Mode [14], Windsurf Cascade [15], Cline [40], and Amazon Q Developer [41].
Our focus was on IDE-based agents, as they offer more flexibility for developers to choose whether to rely fully on the agent or closely collaborate on tasks. Moreover, IDE-based AI tools like GitHub Copilot, Cursor, and Windsurf have gained significant popularity in recent years.
In assessing these tools’ capabilities, we found that Amazon Q was lacking features present in the other tools, such as output explanations, where an agent provides a textual description of the changes it made and the logic behind them. VSCode Agent Mode was newly released at the time of our study and not yet as mature as other options. Although Cline is a capable SWE agent, the grey literature suggests that it has a ‘steeper learning curve, less polished UI than Cursor or Windsurf’, [^1] https://medium.com/@pahwar/cline-vs-windsurf-vs-pearai-vs-cursor-2025s-top-ai-coding-assistants-compared-2b04b985df51 making it less appropriate for a study with limited time for participants to learn the tool.
This left Cursor Agent and Windsurf Cascade as our final candidates. We conducted pilot testing using both tools with a small convenience sample of developers and found the tools comparable in capabilities and usability. However, Cursor Agent offered one feature that made it particularly suitable for our study: automatic incorporation of users’ currently open files and cursor position as context for the agent. Therefore, we selected Cursor Agent as our study tool. Participants used Cursor v 0.47 with Claude 3.5 Sonnet.
@@ III-B Tasks
  • actively maintained (i.e., had commits in the prior month);
@@
Additionally, we selected repositories that were part of two major GitHub organizations that are managed by our company. This enabled us to select participants based on internal-only information about repository contributors, like seniority and geographic location, and to communicate more easily with those contributors.
While most participants selected their first task to complete based on the list of open issues in their project, some participants selected tasks to work on ahead of time. We further asked participants to choose issues that could be solved based on code edits, rather than configuration or documentation fixes, to ensure the agent could provide meaningful assistance.
If participants had time to work on multiple tasks during the session, after they completed their first task, we collaborated with them to choose subsequent issues to ensure diversity in:
  • task type (e.g., if the first issue was a bug, we encouraged participants to next select a feature request), and
  • difficulty and scope (e.g., if the first issue was completed easily, we encouraged participants to next select an issue that seemed more challenging).
@@ III-C Participants
After shortlisting repositories, we reached out to contributors who were employees of our company and who had contributed at least one commit to the repository in the previous 3 months. Additionally we aimed to recruit a diverse sample of participants by reaching out to participants belonging to different regions, genders, levels of seniority and with different levels of activity in the repository. 26 candidates declined or ignored our invitation to participate. We also recruited two participants based on recommendations from other invitees who were unavailable. In total, 19 participants accepted our invitation and completed the study. Participants demographics are presented in Table I. 18 of our participants reported having experience with Github Copilot Chat in their regular programming work, while 3 participants reported using Cursor.
@@ III-D Protocol
Each study session was a meeting between the participant and the first author, conducted using a video conferencing tool. Sessions were scheduled for 60 minutes, with some variability based on configuration issues and participants’ availability beyond the scheduled time. At the start of each session, the study administrator asked the participant to fill out a pre-study questionnaire and consent form. Following this, the study administrator gave a tutorial of the Cursor IDE and its interface by working through a demonstration task to ensure that participants understood the tool, its usage, and its features. The study administrator instructed participants to think aloud and to focus on solving the issue, rather than assessing the AI, to create a more realistic environment. Participants were also instructed to try to use the AI for all their tasks, and make small manual changes if necessary – this instruction was added after some pilot study participants stopped using the agent altogether after switching to manual editing.
@@
Participants went through the list of open issues of the repository on GitHub to choose one to work on. [^2] Although we told participants that they did not need to prepare for the study, five participants planned to address a specific work item during the study. We allowed them to continue as they had planned. Once they completed an issue, the study administrator first asked the participant if they were satisfied enough with the fix to submit it as a pull request. If not, we asked them to keep working to get to that level of satisfaction (if possible within the scope of the session) before moving on to the next issue. We did not ask participants to create and submit a pull request since that would require them to enter their GitHub credentials onto the study administrator’s system. Nonetheless, so that participants could create a pull request with their credentials, we sent them the diff of their work as a patch after the session.
@@ III-E Data Collection
TABLE II: Participant Activity Taxonomy
Activity Description
IDE Actions
Checkpoint Revert to an earlier checkpoint in the conversation
Accepted Code Change Accept code changes proposed by the agent
Rejected Code Change Reject code changes proposed by the agent
Stopped Agent Stop agent run prematurely
Communication
Prompt Crafting Writing a prompt in the chat window
Prompt Submitted Submitting a prompt in the chat window
Following Agent Live Waiting for/ following the steps of agent while it runs
Thinking/Verifying Explanations Actively thinking about and/or verifying textual information by the agent
Thinking/Verifying Actions Actively thinking about the actions taken by the agent to reach its final output (e.g. searching codebase for relevant files)
Thinking/Verifying Code Actively thinking about and/or verifying a proposed or accepted code change by the agent
Code Actions
Writing New Code Manually writing code that implements new functionality
Debugging/Testing Code Manually running or debugging code to check functionality (e.g. testing UI changes on local browser)
Editing Suggested Code Manually editing code changes suggested by the agent
Editing (Personally) Written Code Editing code written by a programmer that is not a suggestion by the agent for the purpose of fixing existing functionality
Non-Code Actions
Looking up issue Referring to issue description/ details
Looking up Documentation Checking an external source for the purpose of understanding code functionality or other information needed to help work on the issue (e.g. Stack Overflow)
Waiting for Terminal Waiting for terminal executions to complete
Deferring Thought For Later Programmer accepts suggestion without verifying/understanding it, but plans to verify it after
Thinking About New Code To Write Thinking about what code or functionality to implement and write (e.g. reading/navigating through the existing codebase)
TABLE III: Chat Trajectory Taxonomy
Code Description
Information Given to Agent in Prompt
Issue Knowledge Providing agent with instructions/knowledge about the issue based solely on original issue description, if any
Implementation Knowledge Providing agent with instructions/knowledge about the issue related to implementation/configuration/scope etc., not solely based on original issue description or environmental context
Styling Knowledge Providing agent with instructions/knowledge about the issue related to code styling/practices not solely based on original issue description or environmental context
Affirmation Affirming the follow up provided by the agent
Meta-details Informing the agent of plan-level heuristics or steps to follow in its response, not grounded in the codebase
Response Sought from Agent in Prompt
Change Explanation Seeking explanations of code changes made by the agent
Codebase Explanation Seeking explanations pertaining to the existing codebase
External Explanation Seeking explanations pertaining to external knowledge not rooted solely in the codebase
Refine Change Seeking corrected version of old code change
New Change Seeking new code changes
Testing Seeking to test existing/previously generated code
Agent Actions
New code change Add code changes to a file
Refine code change Refine previously made changes to a file
Run terminal command Run terminal command
Explain plan Textual description of multi-step plan to be performed by the agent
Explain codebase Explains details of existing code in the codebase
Explain old changes Provides textual explanations of code changes made in previous runs
Explain new changes Provides textual explanations of code changes made in current run
Search Web Searches the internet
Explain external Explains external details (usually based on web search)
Search codebase Searches through the codebase
Analyse files Reads/analyses files in the codebase
Follow Up Prompts the participant for follow-up queries (usually at end of response)
@@ III-E1 Session recordings
To understand participants’ usage of the agent in the session, we used video, audio, and screen recordings to qualitatively coded both participant actions as well as the details of their interactions with the agent.
@@
We coded participant actions by adapting the CUPS taxonomy, previously used for understanding common programmer actions when using AI code completion [21]. We modified this taxonomy to add states unique to collaborating with SWE agents, such as following the trajectory of the agent’s execution and backtracking to a previous point in the conversation (Figure 1a). Table II describes the specific codes we applied.
@@
We also coded the trajectories of the developer-agent interactions in our study by coding information on the context the participant provides to the agent (for example, files, code snippets, and external links) as well as the information they give to and seek from the agent in their prompts. To create our codebook, we performed open coding on two participants’ traces, then iterated on this codebook by annotating trajectories for two further participants to create the final version of our codebook, as described in Table III.

### v3

# Why AI Agents Still Need You: Findings from Developer-Agent Collaborations in the Wild

@@ Abstract
Software Engineering Agents (SWE agents) can autonomously perform development tasks on benchmarks like SWE Bench, but still face challenges when tackling complex and ambiguous real-world tasks. Consequently, SWE agents are often designed to allow interactivity with developers, enabling collaborative problem-solving. To understand how developers collaborate with SWE agents and the barriers they face in such interactions, we observed 19 developers using an in-IDE agent to resolve 33 open issues in repositories to which they had previously contributed. Participants successfully resolved about half of these issues, with those solving issues incrementally having greater success than those using a one-shot approach. Participants who actively collaborated with the agent and iterated on its outputs were also more successful, though they faced challenges in trusting the agent’s responses and collaborating on debugging and testing. Our findings suggest that to facilitate successful collaborations, both SWE agents and developers should actively contribute to tasks throughout all stages of the software development process. SWE agents can enable this by challenging and engaging in discussions with developers, rather than being conclusive or sycophantic.
@@ I Introduction
SWE agents have shown promising results on established benchmarks of software development tasks, such as fully autonomously solving many of the GitHub issues in SWE-Bench [3]. However, they still struggle to solve certain issues due to factors such as the complexity and ambiguity of the issues, lack of tacit knowledge, and limited access to the full development environment [4, 5]. To overcome these limitations, many SWE agents are designed to interact with developers, enabling a collaborative approach to problem-solving. This interaction allows developers and agents to share responsibilities such as understanding the codebase, localising, coding, debugging, testing, and reviewing each other’s work.
Figure 1: Example usage of Cursor Agent (not from our study)
Figure 1 shows an example of one such SWE agent – Cursor Agent in the Cursor integrated development environment (IDE) – responding to a developer’s prompt that has been copied from a GitHub bug report [6] (this specific bug was chosen purely to illustrate the usage of in-IDE SWE agents and was not part of our study). In the top part of the figure, the agent searches the codebase, describes the issue based on this search, and tries to implement a fix. After prompting, the agent then adds a unit test and runs it. While reviewing the agent’s changes, the developer notices a functional call erroneously changed by the agent (Figure 2, bottom), and rejects that part of the change, before reviewing the remaining 5 changes.
There is reason to believe that collaborations between agents and developers will not be smooth and painless. For non-agentic tools, prior research has found that developers sometimes struggle with understanding and trusting the AI’s output, hindering effective collaboration [7, 8, 9]. Is this also the case for SWE agents? On one hand, unlike non-agentic tools that generate a single response and thus act as a black box, agents can provide developers with information about their process by showing their work as they go. Further, the capability of agents to make decisions autonomously allows for developers to express higher-level goals rather than low-level instructions. On the other hand, the broader scope of agents’ decisions and subsequent actions can exacerbate trust issues and place a higher cognitive load on the developer of understanding both the agent’s reasoning trajectory as well as the mutations it performs to the codebase. The answer is yet unclear, due to a lack of research on real-world usage of SWE agents [2].
@@ 2. RQ2:
    What barriers do developers face when collaborating with SWE agents to resolve open issues in active codebases?
@@ 3. RQ3:
    What factors influence participants’ success when collaborating with SWE agents to resolve open issues in active codebases?
@@
Figure 2: A suggested code edit diff by the agent.
@@ II-A Agentic Tools for Software Development
In software development, there has been a rapid rise in the number of agentic tools in research and in practice [10]. Many of these tools aim to solve end-to-end software engineering tasks, as reflected in the evolution of benchmarks such as SWE-bench [3], a set of real previously-solved GitHub issues, and LiveCodeBench [13], which adds new issues over time. Similarly, SWE-Lancer [4] uses real-world freelance software engineering tasks to evaluate AI tools. While these benchmarks aim to evaluate real-world performance, they all test the autonomous performance of agentic tools, rather than their performance with a human-in-the-loop. As Liu and colleagues report, most research-based agentic systems target maximum automation, where human involvement is limited to specification of the problem statement [2].
@@ II-B Developer Experiences with AI Tools
Prior research has explored the strategies that developers apply when using AI [17], especially those related to prompt engineering [18, 19, 20]. Prior work has also found that several development tasks are time-consuming or challenging when working with AI developer tools, including communicating the required context to AI tools [8, 17], verifying AI-generated code recommendations [21], and debugging incorrect AI-generated code [9]. Developer studies are mixed about whether AI increases developer efficiency [9] or decreases it overall [22]. Studies have also found that developers often use AI tools as learning aids and information sources [9, 23], and have concerns about the security [9, 22], quality and reliability [21, 7] of generated code. While these studies have richly explored developer interactions with AI code completion [21, 9, 8, 24] and AI-based chat [25, 26, 27, 28], there is a dearth of developer experience studies on the use of SWE agents. Yet, SWE agents offer a unique interaction paradigm, differing from non-agentic tools in their capability to perform complex tasks autonomously and take feedback from the environment [2].
@@
II-C Communication in Software Development Teams
Prior work has found that effective communication of knowledge in software engineering is a complex problem that requires active involvement from all collaborators. Rus and colleagues report that the majority of software engineering knowledge is tacit, that is, gained through personal experience, rather than explicitly documented. This can make it difficult for new team members to work effectively, as they lack the knowledge that current team members have [5]. Gonçalves and colleagues correspondingly reported that developers spent the most amount of time in collaborative work if they were a new team member or if they had business/customer knowledge [35]. Prior literature further suggests that a lack in streamlined, synchronous communication can result in mistrust and frustration among team members, particularly in distributed software development teams [36, 37]. It is thus important for both managers [38] as well as junior software developers [39] to be effective communicators and collaborators.
Kuttal and colleagues [40, 41] found through Wizard-of-Oz studies that developer-agent interactions tend to suffer due to a lack of extended discussions and the agent not understanding non-verbal cues. Extending this line of work, we aim to leverage findings on communication barriers in software development teams to investigate such barriers in collaborations between developers and in-IDE SWE Agents.
@@ III Study Design
To answer our research questions, we aimed to observe developers collaborating with a SWE agent in a highly ecologically valid setting (in ‘the wild’). To achieve this, we asked 19 professional developers to use Cursor Agent to resolve issues in a codebase to which they have previously contributed.
@@ III-A Tool Selection
In selecting the appropriate SWE agent for our study, we considered the spectrum of human-AI collaboration capabilities, ranging from highly autonomous agents to those that enable flexible interactions with developers. We reviewed several agents, starting with fully autonomous options like AutoCodeRover [42], and moving to agents that allow a moderate level of human-AI collaboration. These moderate collaboration tools often operate outside the IDE and primarily rely on user interaction through prompting, such as Devin [43], Aider [44],OpenHands [45] and GitHub Copilot Workspaces [46]. Finally, we considered agents that operate within an IDE, facilitating high collaboration not only through prompting but also by providing developers access to the full development environment. Examples include Cursor Agent [16], VSCode Agent Mode [14], Windsurf Cascade [15], Cline [47], and Amazon Q Developer [48].
We decided to focus on IDE-based agents as they offer more flexibility for developers to choose whether to rely fully on the agent or closely collaborate on tasks. Moreover, IDE-based AI tools like GitHub Copilot, Cursor, and Windsurf have gained significant popularity in recent years.
To select among these candidate tools, we assessed the types of features and interaction mechanisms present in each of these tools (as of April 30, 2025). These features included those related to flexibility in interactions (modelessness, ability to preview code changes, ability to backtrack within conversations, ability to synchronously edit the same file as the agent), context used by the agent (files opened in IDE, recent user actions) and the interpretability of the agent’s outputs (code change explanations, visualization of agent’s reasoning). We found that Amazon Q was lacking some of these features, such as providing explanations of its outputs. VSCode Agent Mode was newly released at the time of our study and not yet as mature as other options. Although Cline is a capable SWE agent, the grey literature suggests that it has a ‘steeper learning curve, less polished UI than Cursor or Windsurf’ [49], making it less appropriate for a study with limited time for participants to learn the tool.
This left Cursor Agent and Windsurf Cascade as our final candidates. Both of these tools contained almost all of the features that we found in our candidates. We conducted pilot testing using both tools with a small convenience sample of developers and found the tools comparable in capabilities and usability. However, Cursor Agent offered one feature that made it particularly suitable for our study: automatic incorporation of users’ currently open files and cursor position as context for the agent. Therefore, we selected Cursor Agent as our study tool. Participants used Cursor v0.47 with Claude 3.5 Sonnet.
@@ III-B Tasks
  • actively maintained (i.e., commits in the prior month);
@@
Further, we limited our selection to open-source repositories that were part of two major GitHub organizations that are managed by our company. This enabled us to select participants based on internal-only information about repository contributors, like seniority and geographic location, and to communicate more easily with those contributors.
While most participants selected their first task based on the list of open issues in their project, some (five) participants selected tasks to work on before the start of the session, some of which were not GitHub issues. We allowed them to continue as they had planned. We further asked participants to choose issues that could be solved based on code edits, rather than configuration or documentation fixes, to ensure the agent could provide meaningful assistance.
If participants had time to work on multiple tasks during the session, after they completed their first task, we collaborated with them to choose subsequent issues to ensure diversity in task type (e.g., if the first issue was a bug, we encouraged participants to next select a feature request) and difficulty (e.g., if the first issue was completed easily, we encouraged participants to next select an issue that seemed more challenging).
@@ III-C Participants
After shortlisting repositories, we reached out to a total of 43 contributors who were employees of our company and who had contributed at least one commit to the repository in the previous 4 months. We aimed to recruit a diverse sample of participants by reaching out to participants belonging to different regions, genders, levels of seniority and with different levels of activity in the repository. 26 candidates declined or ignored our invitation to participate. We recruited two participants based on recommendations from other invitees who were unavailable. In total, 19 participants accepted our invitation and completed the study. Participants demographics are presented in Table I. 18 of our participants reported having experience with Github Copilot Chat in their regular programming work, while 3 participants reported using Cursor.
@@ III-D Protocol
Each study session was a meeting between the participant and the first author, conducted using a video conferencing tool. Sessions were scheduled for 60 minutes, with some variability based on configuration issues and participants’ availability beyond the scheduled time. To ensure that such variations would not affect participants’ behavior, we informed them that the objective of the study was to observe their interactions with the agent rather than to solve as many issues as possible. Further, we did not analyze time-dependent variables such as action durations and counts, focusing instead on the presence or absence of these actions (as described in Section III-F).
At the start of each session, the study administrator asked the participant to fill out a pre-study questionnaire and consent form. Following this, the study administrator gave a tutorial of the Cursor IDE and its interface by working through a demonstration task to ensure that participants understood the tool, its usage, and its features. Participants were instructed to think aloud and to focus on solving the issue, rather than assessing the AI, to create a more realistic environment. They were also instructed to try to use the AI for all their tasks, and make small manual changes if necessary – this instruction was added after some pilot study participants stopped using the agent altogether after switching to manual editing.
@@
Once participants completed an issue, the study administrator first asked the participant if they were satisfied enough with the fix to submit it as a pull request. If not, participants were asked to keep working to reach that level of satisfaction (if possible within the scope of the session) before moving on to the next issue. We did not ask participants to create and submit a pull request since that would require them to enter their GitHub credentials onto the study administrator’s system. Nonetheless, so that participants could create a pull request with their credentials, we sent them the diff of their work as a patch after the session.
@@ III-E1 Session recordings
To understand participants’ usage of the agent in the session, we used video, audio, and screen recordings to qualitatively code both participant actions as well as the details of their interactions with the agent.
@@
We coded participant actions by adapting the CUPS taxonomy, previously used for understanding common programmer actions when using AI code completion [21]. We modified this taxonomy to add states unique to collaborating with SWE agents, such as following the trajectory of the agent’s execution and backtracking to a previous point in the conversation (Figure 1a).
@@
We also examined the trajectories of the developer-agent interactions in our study by coding information on the context the participant provides to the agent (e.g., files, code snippets, and external links) as well as the information they gave to and sought from the agent in their prompts. To create our codebook, we performed open coding on two participants’ traces, then iterated on this codebook by annotating trajectories for two further participants to create the final version of our codebook.